OpenStudy (anonymous):

11. Your realized income is $3,167.89/month, and your fixed expenses are $954.32/every 2 weeks. If you save 50% of your discretionary monies each month, how much are you saving? (1 point)

OpenStudy (anonymous):

Take 3,167.89 - 954. 32 = X - 954.32 = X. Then take X and divide it by 2 to get your answer.

OpenStudy (anonymous):

Thanks (: